OpenStudy (anonymous):

get medals A grocer wants to make a 10-pound mixture of peanuts and cashews that he can sell for $4.75 per pound. If peanuts cost $4.00 per pound and cashews cost $6.50 per pound, how many pounds of cashews should he use? 3 4 6 7

OpenStudy (anonymous):

Let \(c\) and \(p\) denote the pounds of cashews and peanuts, respectively. Then you have the system \[\begin{cases} c+p=10&\text{because you want a 10 pound mix}\\ 4p+6.5c=4.75(c+p)&\text{because you want the mix's price to be }$4.75 \end{cases}\] Solve for \(c\). Should be easy by substitution.